4) Fab Manufacturing Corporation manufactures and sells stainless steel coffee mugs. Expected mug sales at Fab (in units) for the next three months are as follows:

Fab likes to maintain a finished goods inventory equal to 30% of the next month's estimated sales. How many mugs should Fab plan on producing during the month of November?

  23,200 mugs
  26,800 mugs
  25,900 mugs
  34,300 mugs
Answer #1
Concepts and reason

Job order costing: Job order costing is a method of cost accounting, in which cost is collected and accumulated for each job, work order, or project separately. Customized goods produced firms are following job order costing method.

Assets: Asset is the resource used by the company to generate income. Assets can be classified into different types, Current assets, long term assets and intangible assets. Assets that are expected to be converted into cash within one year are called as current assets. Assets that are expected to be converted into cash more than one year are called as long-term assets. Asset which cannot be touch and see is called as intangible asset. Goodwill, patents etc are the examples of intangible assets.

Inventory: Inventory is one of the important current assets of a company. Inventory includes raw materials, work in process and finished goods.

Raw materials: Raw materials inventory is a component used by the company to produce finished goods. Raw material is the basic input used for the production function. Metals, woods, plastics etc are the examples of raw material inventory.

Work-in process: It refers to units that could not convert into the finished goods in the current period. Work-in process is also known as semi manufactured products.

Finished goods: Finished goods are the completely finished product ready for sale.

Fundamentals

Beginning inventory: Beginning inventory is the inventory at the beginning period. Last years’ ending inventory and current years’ beginning inventory are same.

Ending inventory: Ending inventory is the remaining inventory at the end of the period. Some company’s desired ending inventory is certain percentage of the next month's expected unit sales. Current years’ ending inventory and next years’ beginning inventory are same.

Production units: Production units are the total units produced by the company. It can be calculated by subtracting the beginning inventory from the sum of ending inventory and expected sales. The following formula is used for calculating the production units.

Productionunits=(Endinginventory+ExpectedsalesBeginninginventory){\rm{Production units = }}\left( \begin{array}{l}\\{\rm{Ending inventory }}\\\\{\rm{ + Expected sales }}\\\\ - {\rm{ Beginning inventory}}\\\end{array} \right)

Expected sales: An expected sale is the unit of inventory a company expects to sale at some point in the future.

Calculation of ending inventory in November is given below:

EndinginventoryinNovember=(Decembermonthsalesunit×Percentageofnextmonthsales)=31,000mugs×30%=9,300mugs\begin{array}{c}\\{\rm{Ending inventory in November = }}\left( \begin{array}{l}\\{\rm{December month sales unit }}\\\\{\rm{ \times Percentage of next month sales}}\\\end{array} \right)\\\\{\rm{ = 31,000 mugs \times 30\% }}\\\\{\rm{ = 9,300 mugs}}\\\end{array}

Therefore, the ending inventory in November is 9,300 mugs.

Calculation of the F Manufacturing Corporation's production units in November is given below:

ProductionunitsinNovember=(EndinginventoryinNovember+SalesinNovemberBeginninginventoryinNovember)=9,300mugs+25,000mugs7,500mugs=26,800mugs\begin{array}{c}\\{\rm{Production units in November = }}\left( \begin{array}{l}\\{\rm{Ending inventory in November}}\\\\{\rm{ + Sales in November}}\\\\ - {\rm{ Beginning inventory in November}}\\\end{array} \right)\\\\{\rm{ = 9,300 mugs + 25,000 mugs }} - {\rm{ 7,500 mugs}}\\\\{\rm{ = 26,800 mugs}}\\\end{array}

Therefore, the production units in November are 26,800 mugs.

Working note:

Calculation of beginning inventory in November is given below:

BeginninginventoryinNovember=(Novembermonthsalesunit×Percentageofnextmonthsales)=25,000mugs×30%=7,500mugs\begin{array}{c}\\{\rm{Beginning inventory in November = }}\left( \begin{array}{l}\\{\rm{November month sales unit }}\\\\{\rm{ \times Percentage of next month sales}}\\\end{array} \right)\\\\{\rm{ = 25,000 mugs \times 30\% }}\\\\{\rm{ = 7,500 mugs}}\\\end{array}

Therefore, the beginning inventory in November is 7,500 mugs.

Ans:

Production units in November are 26,800 mugs.